solve the following equations and express the answer as an ordered pair if possible
2x-4y=8
x-2y=4

Answer by funmath(2933)

solve the following equations and express the answer as an ordered pair if possible
L1) 2x-4y=8
L2) x-2y=4
Multiply L2 by -2 and add it to L1 in order to eliminate one of the variables.
-2(x-2y=4) --->> -2x+4y=-8
:
2x-4y=8
-2x+4y-8
___________
0x+0y=0
0=0
:
This is a special case, our variables are both eliminated and our constants are =. Graphically, they are the same line and have all points in common.
This means there are an infinite number of solutions. Some teachers and books are happy to get that answer or "all real numbers."
:
Others want you to solve either equation for y and write your answer like this:
x-2y=4
-2y=-x+4
-2y/-2=-x/-2+4/-2
y=(1/2)x-2
The answer is: (x,(1/2)x-2)
You don't have to do this last part unless your teacher or book requires it.
:
Happy Calculating!!!